# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

### v4

@@ Abstract
Despite the rapid growth of machine learning research, corresponding code implementations are often unavailable, making it slow and labor-intensive for researchers to reproduce results and build upon prior work. In the meantime, recent Large Language Models (LLMs) excel at understanding scientific documents and generating high-quality code. Inspired by this, we introduce PaperCoder, a multi-agent LLM framework that transforms machine learning papers into functional code repositories. PaperCoder operates in three stages: planning, where it constructs a high-level roadmap, designs the system architecture with diagrams, identifies file dependencies, and generates configuration files; analysis, which focuses on interpreting implementation-specific details; and generation, where modular, dependency-aware code is produced. Moreover, each phase is instantiated through a set of specialized agents designed to collaborate effectively across the pipeline. We then evaluate PaperCoder on generating code implementations from machine learning papers based on both model-based and human evaluations, particularly from the authors of those papers, with author-released repositories as ground truth if available. Our results demonstrate the effectiveness of PaperCoder in creating high-quality, faithful implementations. Furthermore, it consistently shows strengths in the recently released PaperBench benchmark, surpassing strong baselines by substantial margins. Code is available at: https://github.com/going-doer/Paper2Code.
@@ 2 Related Work
LLMs have shown impressive capabilities in text understanding and generation (OpenAI, 2024; Dubey et al., 2024; Reid et al., 2024) and widely utilized for specialized domains (beyond general tasks), such as mathematics, science, and coding (Prabhakar et al., 2025; Wang et al., 2024; Trinh et al., 2024). Particularly, code-specialized LLMs (Hui et al., 2024; DeepSeek-AI et al., 2024; 2025) have received significant attention thanks to remarkable performance on various software engineering tasks (Xia et al., 2024), including software design and development (Qian et al., 2024; Hong et al., 2024), requirements elicitation (Mu et al., 2023), and formal specification generation (Luo et al., 2024). Our work aligns closely with this line of research, exploring and expanding upon the capabilities and applications of (code-specialized) LLMs.
@@ 4.1 Experimental Setup
To evaluate our PaperCoder, we construct a new benchmark (Paper2CodeBench). Specifically, we collect the accepted papers from recent machine learning venues (such as ICLR, ICML, and NeurIPS 2024) with the OpenReview API [^1] https://docs.openreview.net/reference/api-v2, and filter them based on the availability of code with its total number of tokens less than 70,000, to ensure the full repository remains within reasonable processing limits of modern LLMs for generation and evaluation. Also, to maintain the quality, we perform model-based evaluation (Liu et al., 2023) with GPT-4o on all the collected repositories and select the top 30 from each venue, resulting in a total of 90 papers listed in Tables 1718, and 19. Moreover, we additionally consider 21 papers for human evaluation (See Table 20). In addition to Paper2CodeBench, we also use the recently released PaperBench Code-Dev (Starace et al., 2025), which consists of 20 papers from ICML 2024 with paper-specific rubrics annotated by humans. In particular, those rubrics are used to judge the correct implementation based on LLM-based evaluation.
@@ 4.2 Experimental Results and Analysis
To see how much each component of PaperCoder contributes to the performance gain, we conduct ablation studies on the subset of Paper2CodeBench (composed of ICML papers). Specifically, we start with the method that uses only the full paper and incrementally add components in the order they are executed (such as overall plan, architecture design, logic design, configuration generation, and final analysis), reported in Table 6. We then observe that the performance steadily improves as additional components are incorporated. Meanwhile, a performance drop occurs when the architecture design module is added; however, while this might seem surprising at first, it is in fact expected: architecture design alone does not specify the execution or implementation order of files, which leads to confusion during the code generation stage. However, this issue is addressed once the subsequent logic design module explicitly defines file dependencies and establishes a clear generation order. Overall, integrating all modules in the pipeline yields the highest performance, confirming the effectiveness of our fully structured, multi-stage pipeline with various modules proposed.
@@
We confirm in Table 6 that the planning and analysis stages play a pivotal role in guiding subsequent analysis and coding, and we further test whether refining earlier outputs can improve downstream performance. Specifically, we augment the planning and analysis phases with verification-and-refinement steps (See Figures 19 to 28 for prompts), following Self-Refine (Madaan et al., 2023), and evaluate a total of 30 papers subsampled from Paper2CodeBench (10 from each conference). As shown in Table 7, refinement of planning and analysis improves their own outputs but also leads to measurable gains in the subsequent stages, reducing downstream errors.
@@ 4.3 Additional Analysis on Reproduction from Implemented Code Repository
It is worth noting that making the repository-level code executable and fully reproducible in one go is extremely challenging (even for humans), as demonstrated by Starace et al. (2025). Also, our goal is to provide a faithful starting point that meaningfully aids reproduction efforts (Figure 6), rather than aiming for perfect reproduction. Nevertheless, to assess how close our generated repositories are to being directly executable, we perform manual execution evaluations on five papers. Specifically, when execution fails, we manually debug and refine the code and adapt the input data as needed to enable successful runs. We then find that, on average, only 0.81% of the code lines require minor modification, such as updating deprecated API or correcting data type mismatches, for successful execution (See Examples in Figures 7 and 8 with statistics in Table 13), which highlights that our generated repositories are near-executable with minimal human intervention.
@@ 5 Conclusion
In this work, we introduced PaperCoder, a framework that automatically generates code repositories from research papers in machine learning through a structured, three-stage pipeline. Specifically, we defined a high-level roadmap, system architecture, execution logic, and configuration via the planning stage, which are then enhanced through detailed per-file analysis, followed by the sequential code generation informed by artifacts from prior stages. To validate PaperCoder, we performed evaluations on two benchmarks: our Paper2CodeBench, comprising recent papers from top-tier machine learning venues, and (recently released) PaperBench Code-Dev, providing fine-grained evaluation protocols, on which PaperCoder consistently outperforms existing baselines on both model-based and human evaluations. Furthermore, additional analyses demonstrate its robustness and practicality: it remains effective across different LLM backbones, shows strong executability with only 0.81% of the lines requiring minor fixes, and benefits from each stage in the pipeline. We envision PaperCoder as one important step toward accelerating scientific progress by aiding the reproduction of research papers.
@@
We attach the code to reproduce our work in the supplementary materials. Detailed instructions for running the experiments are included in the accompanying README files, and furthermore, all necessary details to reproduce our experiments are described in Section 4.1 and in Appendix A.1.
@@ B.1 Code Availability
B.2 PaperBench Code-Dev Results
@@
B.3 Impact of Paper Content on Code Generation
@@
Ref-based
Average Score
4.26 (0.28) 3.75 (0.55)
@@
B.4 Most Common Types of Errors and Failure Modes
@@
B.5 Analysis of Performance Across Paper Categories
@@ Appendix C Limitations and Future Work
While PaperCoder demonstrates strong performance in reproducing machine learning papers (where code implementations are particularly helpful and usually necessary for validating research ideas), its current scope is limited to this domain. Beyond this, we believe accelerating the reproduction of scientific discovery to other domains where code is not the primary medium for validation, such as theoretical mathematics, is an exciting direction for future work. In addition, the current version of PaperCoder processes only textual inputs, and extending it to process visual inputs (such as figures in papers) is an interesting avenue. Lastly, as with other repository-level code generation approaches, improving executability remains an important (but still challenging) direction for future work.
@@
Table 14: Qualitative analysis of top-ranked repositories. We categorize the reasons why human annotators select the repositories generated by our PaperCoder framework as their top choice into six (described in the first row).
@@
Repo Name
Analysis of Reproducibility
CoLoR
Execution was successful, but the ORPO loss was likely mis-specified, causing the compression model to fail in training as intended. This issue stems from the overly simplified description of the loss function in the paper.
cognitive-behaviors
Successfully reproduced SFT and RL training processes but encountered a minor error in parsing model responses during evaluation.
RADA
Implementation closely matched the paper, but missing details prevented full reproduction of the reported results, leading to identical samples.
Self-Instruct
Executed smoothly and accurately reflected the procedure described in the paper.
G-EVAL
Implemented only the Coherence metric, though the original paper included Coherence, Consistency, Fluency, and Relevance. The Coherence implementation was faithful and correct.
@@
Original 1# data_loader.py 2import openai 3import pandas as pd 4from rouge_score import rouge_scorer 5 6# Set up basic logging configuration 7logging.basicConfig(level=logging.INFO, format="%(asctime)s [%(levelname)s] %(message)s") 8 9 10...(omited)... 11 12 response = openai.Completion.create( 13 engine=self.engine, 14 prompt=prompt, 15 max_tokens=150, 16 temperature=0.7, 17 n=1, 18 stop=["Task 16"] 19 ) 20 raw_text = response.choices[0].text.strip()
Editied 1# data_loader.py 2from openai import OpenAI 3import pandas as pd 4from rouge_score import rouge_scorer 5 6# Set up basic logging configuration 7logging.basicConfig(level=logging.INFO, format="%(asctime)s [%(levelname)s] %(message)s") 8 9openai_api_key = os.environ.get("OPENAI_API_KEY") 10client = OpenAI(api_key=openai_api_key) 11 12...(omited)... 13 14 response = client.chat.completions.create( 15 model=self.engine, 16 messages=[{"role": "user", "content": prompt}], 17 max_completion_tokens=150, 18 temperature=0.7, 19 n=1, 20 stop=["Task 16"] 21 ) 22 raw_text = response.choices[0].message.content.strip()
Figure 7: Case study on the reproduction of the Self-Instruct paper. The left shows the code generated by PaperCoder using o3-mini-high, and the right shows the version manually edited by the authors to correct the error. In this example, an outdated API call is updated to its latest version. In the initial version, lines 2, 12, 14, 15, and 29 are removed; in the edited version, lines 2, 9, 10, 14, 16, 17, and 22 are added.
Original 1# trainer.py 2self.optimizer = AdamW(self.model.model.parameters(), lr=lr) 3 4# model.py 5self.model = AutoModelForCausalLM.from_pretrained(base_model)
@@
Edited 1# trainer.py 2self.optimizer = AdamW(self.model.model.parameters(), lr=float(lr)) 3 4# model.py 5self.model = AutoModelForCausalLM.from_pretrained(base_model, trust_remote_code=True)
Figure 8: Case study on the reproduction of the CoLoR paper. The left shows the code generated by PaperCoder using o3-mini-high, and the right shows the manually edited version by the authors. In this example, a numeric value is cast correctly, and a required argument is added to enable execution. Lines 2 and 5 are modified.

### v5

# arXiv:2504.17192v5[cs.CL] 27 Feb 2026

@@ Abstract
Despite the rapid growth of machine learning research, corresponding code implementations are often unavailable, making it slow and labor-intensive for researchers to reproduce results and build upon prior work. In the meantime, recent Large Language Models (LLMs) excel at understanding scientific documents and generating high-quality code. Inspired by this, we introduce PaperCoder, a multi-agent LLM framework that transforms machine learning papers into operational code repositories. PaperCoder operates in three stages: planning, where it constructs a high-level roadmap, designs the system architecture with diagrams, identifies file dependencies, and generates configuration files; analysis, which focuses on interpreting implementation-specific details; and generation, where modular, dependency-aware code is produced. Moreover, each phase is instantiated through a set of specialized agents designed to collaborate effectively across the pipeline. We then evaluate PaperCoder on generating code implementations from machine learning papers based on both model-based and human evaluations, particularly from the authors of those papers, with author-released repositories as ground truth if available. Our results demonstrate the effectiveness of PaperCoder in creating high-quality, faithful implementations. Furthermore, it consistently shows strengths in the recently released PaperBench benchmark, surpassing strong baselines by substantial margins. Code is available at: https://github.com/going-doer/Paper2Code.
@@ 2 Related Work
LLMs have shown impressive capabilities in text understanding and generation (OpenAI, 2024; Dubey et al., 2024; Reid et al., 2024) and widely utilized for specialized domains (beyond general tasks), such as mathematics, science, and coding (Prabhakar et al., 2025; Wang et al., 2024b; Trinh et al., 2024). Particularly, code-specialized LLMs (Hui et al., 2024; DeepSeek-AI et al., 2024; 2025) have received significant attention thanks to remarkable performance on various software engineering tasks (Xia et al., 2024), including software design and development (Qian et al., 2024; Hong et al., 2024), requirements elicitation (Mu et al., 2023), and formal specification generation (Luo et al., 2024). Our work aligns closely with this line of research, exploring and expanding upon the capabilities and applications of (code-specialized) LLMs.
@@ 4.1 Experimental Setup
To evaluate our PaperCoder, we construct a new benchmark (Paper2CodeBench). Specifically, we collect the accepted papers from recent machine learning venues (such as ICLR, ICML, and NeurIPS 2024) with the OpenReview API [^1] https://docs.openreview.net/reference/api-v2, and filter them based on the availability of code with its total number of tokens less than 70,000, to ensure the full repository remains within reasonable processing limits of modern LLMs for generation and evaluation. Also, to maintain the quality, we perform model-based evaluation (Liu et al., 2023) with GPT-4o on all the collected repositories and select the top 30 from each venue, resulting in a total of 90 papers listed in Tables 2021, and 22. Moreover, we additionally consider 21 papers for human evaluation (See Table 23). In addition to Paper2CodeBench, we also use the recently released PaperBench Code-Dev (Starace et al., 2025), which consists of 20 papers from ICML 2024 with paper-specific rubrics annotated by humans. In particular, those rubrics are used to judge the correct implementation based on LLM-based evaluation.
@@ 4.2 Experimental Results and Analysis
To see how much each component of PaperCoder contributes to the performance gain, we conduct ablation studies on the subset of Paper2CodeBench (composed of ICML papers). Specifically, we start with the method that uses only the full paper and incrementally add components in the order they are executed (such as overall plan, architecture design, logic design, configuration generation, and analysis), reported in Table 6. From this, we observe that the performance steadily improves as additional components are incorporated. Meanwhile, a performance drop occurs when the architecture design module is added; however, while this might seem surprising at first, it is in fact expected: architecture design alone does not specify the execution or implementation order of files, which leads to confusion during the code generation (see Figure 7). However, this issue is addressed once the subsequent logic design module explicitly defines file dependencies and establishes a clear generation order. Overall, integrating all modules yields the highest performance, confirming the effectiveness of our fully structured, multi-stage pipeline with various modules proposed.
@@
We confirm in Table 6 that the planning and analysis stages play a pivotal role in guiding subsequent analysis and coding, and we further test whether refining earlier outputs can improve downstream performance. Specifically, we augment the planning and analysis phases with verification-and-refinement steps (See Figures 23 to 32 for prompts), following Self-Refine (Madaan et al., 2023), and evaluate a total of 30 papers subsampled from Paper2CodeBench (10 from each conference). As shown in Table 7, refinement of planning and analysis improves their own outputs but also leads to measurable gains in the subsequent stages, reducing downstream errors.
@@ 4.3 Additional Analysis on Reproduction from Implemented Code Repository
It is worth noting that making the repository-level code executable and fully reproducible in one go is extremely challenging (even for humans), as demonstrated by Starace et al. (2025). Also, our goal is to provide a faithful starting point that meaningfully aids reproduction efforts (Figure 6), rather than aiming for perfect reproduction. Nevertheless, to assess how close our generated repositories are to being directly executable, we perform manual execution evaluations on five papers. Specifically, when execution fails, we manually debug and refine the code and adapt the input data as needed to enable successful runs. We then find that, on average, only 0.81% of the code lines require minor modification, such as updating deprecated API or correcting data type mismatches, for successful execution (see examples in Figures 8 to 12 with statistics in Table 15), which highlights that our generated repositories are near-executable with minimal human intervention.
@@ 5 Conclusion
In this work, we introduced PaperCoder, a framework that automatically generates code repositories from research papers in machine learning through a structured, three-stage pipeline. Specifically, we defined a high-level roadmap, system architecture, execution logic, and configuration via the planning stage, which are then enhanced through detailed per-file analysis, followed by the sequential code generation informed by artifacts from prior stages. To validate PaperCoder, we performed evaluations on two benchmarks: our Paper2CodeBench, comprising recent papers from top-tier machine learning venues, and (recently released) PaperBench Code-Dev, providing fine-grained evaluation protocols, on which PaperCoder consistently outperforms existing baselines on both model-based and human evaluations. Furthermore, additional analyses demonstrate its robustness and practicality: it remains effective across different LLM backbones, shows practical executability with only 0.81% of the lines requiring minor fixes, and benefits from each stage in the pipeline. We envision PaperCoder as one important step toward accelerating scientific progress by aiding the reproduction of research papers.
@@
We provide the public GitHub link to the code to reproduce our work. Detailed instructions for running the experiments are included in the accompanying README files, and furthermore, all necessary details to reproduce our experiments are described in Section 4.1 and in Appendix A.1.
Acknoledgements
This work was supported by the Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00256259), the Institute of Information & Communications Technology Planning & Evaluation (IITP) with a grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research (No. RS-2024-00469482 & RS-2024-00509279), and the InnoCORE program of the Ministry of Science and ICT (No. N10250156).
@@ B.1 Code Availability
B.2 Additional Analysis on Executability
Table 11: Developer time comparison.
Method Developer Time
CoLoR 5
cognitive-behaviors 6
RADA 27
self-instruct 53
geval 22
As discussed in Section 4.3, we observe that the fixes required for execution are overwhelmingly simple, which are minor syntax issues, missing imports, or small adjustments to variable names rather than logic- or architecture-level revisions, as shown in Figure 8 to 12 with statistics in Table 15.
To quantify this more concretely, we further estimate developer time by multiplying the number of modified lines by a difficulty factor (1 - 3):
  • 1: simple syntax or typo fixes, variable renaming, comment adjustments
  • 2: fixes requiring local reasoning (e.g., adjusting conditions or API usage)
  • 3: nontrivial issues requiring deeper debugging (asynchronous or memory-related errors)
The results of this estimation are then reported in the developer time column of Table 11, which demonstrates that correcting the errors does not require a significant amount of time.
In addition to human debugging, we also perform LLM-assisted repair to examine whether the generated code can be fixed automatically without human intervention. We use o4-mini-high and GPT-5 during the repair stage. For each repair round, we record (i) the number of iterations required, and (ii) the error categories identified using GPT-5.1. The results, included in Table 14, then show that this automatic repair strategy resolves all the issues mostly within a small number of iterations and that the errors are primarily syntactic or import-related rather than logic-level.
B.3 PaperBench Code-Dev Results
@@
B.4 Impact of Paper Content on Code Generation
@@
Ref-based Average Score 4.26 (0.28) 3.75 (0.55)
@@
B.5 Most Common Types of Errors and Failure Modes
@@
B.6 Addressing Environment Setup
Our framework focuses on faithfully reconstructing the methodological pipeline, rather than fully automating environment configuration. In practice, environment setup tends to be far easier and more reliable for users to adjust manually, whereas reconstructing the core algorithmic logic from natural language is substantially more challenging and central to our goal. For this reason, PaperCoder prioritizes method-level faithfulness over full automation of environment specification, which is highly paper-specific and often varies across systems. Nevertheless, to further handle this environmental issue, we have conducted an additional experiment on 30 papers (10 from each conference in Paper2CodeBench), augmenting PaperCoder with prompts (shown in Figure 33) explicitly aimed at inferring and repairing environment requirements. Then, in these cases, we do not observe dependency-related failures, suggesting that dependency issues in Table 13 are not from the intrinsic weaknesses of the framework. We view incorporating a dedicated environment-construction or “DevOps” agent as an interesting extension for future work.
B.7 Analysis of Performance Across Paper Categories
@@
B.8 Discussion Against Data Contamination in Paper2CodeBench
We evaluate target papers from ICLR/ICML/NeurIPS 2024, because the primary models used in our experiments (OpenAI o3-mini and GPT-4o) have a knowledge cutoff of October 2023, meaning these 2024 paper/code repositories fall after the training window of models. While we cannot fully rule out all forms of indirect exposure, the temporal gap substantially reduces the likelihood of contamination, and thus, data contamination is unlikely to meaningfully affect the results.
B.9 Cross-Family Validation for Paper2CodeBench Model Evaluation
While human evaluation might be the most reliable form of assessment, conducting full-scale human evaluation for all baselines and papers would be prohibitively expensive (Zheng et al., 2023). Therefore, we adopt the standard LLM-as-a-judge strategy for primary assessment. Also, as shown in Table 5, o3-mini-high correlates strongly with human judgments (0.78 in the reference-based setting and 0.73 in the reference-free setting), suggesting that the LLM-based evaluation could be a reliable proxy for human assessments. Nevertheless, to further address the concern on the circularity and bias, specifically that an evaluator from the same model family might favor code generated by the same model family, we perform a cross-family evaluation with Gemini-Flash 2.5 on 30 randomly selected papers of the Paper2Code benchmark. The resulting Spearman correlation of 0.73 between Gemini-Flash 2.5 and o3-mini-high indicates that the evaluation remains consistent across different model families, providing strong evidence that our results are not an artifact of model-family alignment.
@@ Appendix C Limitations and Future Work
While PaperCoder demonstrates strong performance in reproducing machine learning papers (where code implementations are particularly helpful and usually necessary for validating research ideas), its current scope is limited to this domain. Beyond this, we believe accelerating the reproduction of scientific discovery to other domains where code is not the primary medium for validation, such as theoretical mathematics, is an exciting direction for future work. In addition, the current version of PaperCoder processes only textual inputs, and extending it to process visual inputs (such as figures in papers) with recent OCR models capable of extracting figures and tables (in addition to text) (Wang et al., 2024a; He et al., 2024; Wei et al., 2025; Li et al., 2025; Zhang et al., 2025; Niu et al., 2025) is an interesting avenue. Lastly, as with other repository-level code generation approaches, improving executability remains an important (but still challenging) direction for future work.
Table 14: The number of repair iterations (with the LLM) and error categories. We use o4-mini-high and GPT-5 for the automatic repair, and GPT-5.1 for identifying the error categories.
RepoName Iteration Error Category
CoLoR 1 Configuration Update
2 Code Refactor / Robustness Enhancement
3 Bug Fix
4 Performance Adjustment / Behavior Control Update
Cognitive-Behaviors 1 Configuration Update
2 Bug Fix / Type Compatibility Correction
3 Autograd Control / Memory Optimization
4 API Modernization / Correct Mask Handling / Behavior Tuning (Token Budget)
5 Bug Fix / Robustness Improvement
6 Representation Handling Feature + Refactor
7 Configuration Robustness / Type Validation
8 Training Efficiency Optimization / Memory Reduction / Autograd Hygiene
9 Memory & Performance Optimization / Behavioral Correction
10 Configuration Robustness / Type Validation
11 Algorithmic Refactor / Performance & Memory Optimization
API Modernization / Robustness Improvement
RADA 1 Configuration Update
2 API Modernization / Compatibility Update
Self-Instruct 1 Configuration Update
2 API Migration / Compatibility Update
3 API Modernization / SDK Refactor
4 API Rollback / Correct Endpoint Restoration
5 API Consistency Fix
6 API Field Access Modernization
G-EVAL 1 API Modernization / SDK Migration
2 Debug Cleanup / Schema Update
@@
Table 16: Qualitative analysis of top-ranked repositories. We categorize the reasons why human annotators select the repositories generated by our PaperCoder framework as their top choice into six (described in the first row). We also show an example of the response in Table 17.
@@
Table 17: Example responses from human annotators for the reasons for the top-ranked repositories in the human evaluation. We ask the human evaluator (who is the first author of the paper to be reproduced by PaperCoder) the following question: “Among the top-ranked repositories in each group, which one do you think is the best? If the repositories are the same, you may select any of them. Please briefly explain your reasoning.”
RepoName Response for the top-ranked repositories
Janus The code includes everything needed to implement the paper. The code is clean and easy to understand (not overloaded with comments). It’s not overly packaged, and since config files are provided, it’s convenient for running various experiments.
VideoRAG Each component—data processing, retrieval, frame selection, generation, and evaluation—is clearly separated into its own module, making the system easy to maintain and extend. Moreover, the most critical modules are fully implemented, covering the essential functionality required by the framework.
T1 This repository successfully implements the following: Tool-based filtering, Calculate score of each generation after filtering using verifier model, Training code for distillation.
@@
Repo Name Analysis of Reproducibility
CoLoR Execution was successful, but the ORPO loss was likely mis-specified, causing the compression model to fail in training as intended. This issue stems from the overly simplified description of the loss function in the paper.
cognitive-behaviors Successfully reproduced SFT and RL training processes but encountered a minor error in parsing model responses during evaluation.
RADA Implementation closely matched the paper, but missing details prevented full reproduction of the reported results, leading to identical samples.
Self-Instruct Executed smoothly and accurately reflected the procedure described in the paper.
G-EVAL Implemented only the Coherence metric, though the original paper included Coherence, Consistency, Fluency, and Relevance. The Coherence implementation was faithful and correct.
@@
+ Arch. Design 1# main.py 2 3# (... omitted ...) 4 5class DatasetLoader: 6 7# (... omitted ...) 8 9 def load_data(self) -> Tuple[DataLoader, DataLoader, DataLoader]: 10 """Returns DataLoaders for train, validation, and test splits.""" 11 train_dataset = TimeSeriesDataset( 12 dataset_name=self.dataset_name, mode="train", 13 lookback=self.lookback, prediction_horizon=self.prediction_horizon, 14 num_samples=self.num_train 15 ) 16 val_dataset = TimeSeriesDataset( 17 dataset_name=self.dataset_name, mode="val", 18 lookback=self.lookback, prediction_horizon=self.prediction_horizon, 19 num_samples=self.num_val 20 ) 21 test_dataset = TimeSeriesDataset( 22 dataset_name=self.dataset_name, mode="test", 23 lookback=self.lookback, prediction_horizon=self.prediction_horizon, 24 num_samples=self.num_test 25 ) 26 batch_size: int = self.config["training"].get("batch_size", 32) 27 train_loader = DataLoader(train_dataset, batch_size=batch_size, shuffle=True) 28 val_loader = DataLoader(val_dataset, batch_size=batch_size, shuffle=False) 29 test_loader = DataLoader(test_dataset, batch_size=batch_size, shuffle=False) 30 return train_loader, val_loader, test_loader 31 32# (... omitted ...) 33 34class Main: 35 36# (... omitted ...) 37 38 def run_experiment(self) -> None: 39 for dataset_name in dataset_names: 40 for run in range(self.num_runs): 41 loader_instance = DatasetLoader(self.config, dataset_name) 42 train_loader, val_loader, test_loader = loader_instance.load_data() 1# dataset_loader.py 2 3# (... omitted ...) 4 5class DatasetLoader: 6 7# (... omitted ...) 8 9 def load_data(self) -> Tuple[DataLoader, DataLoader, DataLoader]: 10 batch_size: int = int(self.config.get("training", {}).get("batch_size", 32)) 11 train_loader: DataLoader = DataLoader(train_dataset, batch_size=batch_size, shuffle=True, num_workers=0) 12 val_loader: DataLoader = DataLoader(val_dataset, batch_size=batch_size, shuffle=False, num_workers=0) 13 test_loader: DataLoader = DataLoader(test_dataset, batch_size=batch_size, shuffle=False, num_workers=0) 14 return train_loader, val_loader, test_loader
+ Logic Design 1# main.py 2 3# (... omitted ...) 4 5class Main: 6 7# (... omitted ...) 8 9 def run_experiment(self) -> None: 10 num_runs: int = self.config.get("num_runs", 5) 11 seed_base: int = self.config.get("seed", 42) 12 13 for run in range(1, num_runs + 1): 14 dataset_loader = DatasetLoader(self.config) 15 data_loaders = dataset_loader.load_data() 16 train_loader = data_loaders["train"] 17 val_loader = data_loaders["val"] 18 test_loader = data_loaders["test"] 1# dataset_loader.py 2 3# (... omitted ...) 4 5class DatasetLoader: 6 7# (... omitted ...) 8 9 def load_data(self) -> Tuple[DataLoader, DataLoader, DataLoader]: 10 """ 11 Loads the time series data, segments it, splits it into training, validation, and test sets, 12 and returns the corresponding DataLoader objects. 13 14 Returns: 15 Tuple[DataLoader, DataLoader, DataLoader]: (train_loader, val_loader, test_loader) 16 """ 17 # Load CSV using pandas 18 df: pd.DataFrame = pd.read_csv(self.data_file) 19 20# (... omitted ...)
Figure 7: Case study on ablation experiments. Without a dependency-aware generation order, the Architecture Design stage causes main.py to duplicate functions from dataset_loader.py. In contrast, adding the Logic Design phase resolves this issue by aligning the file-generation order.
@@
CoLoR 1# config.yaml 2model: 3 base_model: "microsoft/Phi-3-mini-4k-instruct" 4 alternative_models: 5+ - "mistralai/Mistral-7B-Instruct-v0.3" 6+ - "meta-llama/Llama-3.2-3B-Instruct" 1# trainer.py 2- self.optimizer = AdamW(self.model.model.parameters(), lr=lr) 3+ self.optimizer = AdamW(self.model.model.parameters(), lr=float(lr)) 1# model.py 2- self.model = AutoModelForCausalLM.from_pretrained(base_model) 3+ self.model = AutoModelForCausalLM.from_pretrained( 4+ base_model, trust_remote_code=True 5+ )
Figure 8: Case study on reproducing a paper, called Efficient Long-Context Language Model Retrieval with Compression (Seo et al., 2025) (repository: CoLoR), displaying the files modified during manual debugging.
cognitive-behaviors 1# config.yaml 2- peak_learning_rate: 1e-5 3+ peak_learning_rate: 0.00001 4 5- actor_learning_rate: 1e-6 6- critic_learning_rate: 1e-5 7+ actor_learning_rate: 0.000001 8+ critic_learning_rate: 0.00001 9 10model: 11 base_models: 12- - "Llama-3.2-3B" 13- - "Qwen-2.5-3B" 14+ - "meta-llama/Llama-3.2-3B" 15+ - "Qwen/Qwen2.5-1.5B" 16 17 additional_models: 18- - "Llama-3.1-70B" 19+ - "meta-llama/Llama-3.1-3B"
Figure 9: Case study on reproducing a paper, called Cognitive Behaviors that Enable Self-Improving Reasoners, or, Four Habits of Highly Effective STaRs (Gandhi et al., 2025) (repository: cognitive-behavior), displaying the files modified during manual debugging.
RADA 1# config.yaml 2 t5_base: 3- model_name: "t5-base" 4+ model_name: "google-t5/t5-base" 5 6 llama2_7b: 7- model_name: "Llama2-7B" 8+ model_name: "meta-llama/Llama-2-7b" 9 10- augmentation_model: "Llama2-7B-Chat" 11+ augmentation_model: "meta-llama/Llama-2-7b-chat" 12 13retrieval: 14- embedding_model: "distilbert-base-nli-stsb-mean-tokens" 15+ embedding_model: "sentence-transformers/distilbert-base-nli-mean-tokens" 16 17+ seed_data_path: data/seed_data.json 18+ external_data_path: data/external_data.csv 1# main.py 2- generator: Generator = Generator(llm_model=augmentation_model) 3+ generator: Generator = Generator( 4+ llm_model=augmentation_model, 5+ generation_params={ 6+ "max_length": 2048, 7+ "temperature": 0.7, 8+ "top_k": 50, 9+ "top_p": 0.95, 10+ "num_return_sequences": 1 11+ } 12+ )
Figure 10: Case study on reproducing a paper, called Retrieval-Augmented Data Augmentation for Low-Resource Domain Tasks (Seo et al., 2024) (repository: RADA), displaying the files modified during manual debugging.
self-instruct 1# config.yaml 2- engine: davinci 3+ engine: gpt-4.1-nano 1# dataset_loader.py 2- import openai 3+ from openai import OpenAI 4+ client = OpenAI(api_key=os.environ.get("OPENAI_API_KEY")) 5 6- response = openai.Completion.create( 7- engine=self.engine, 8- prompt=prompt, 9- max_tokens=150, 10+ response = client.chat.completions.create( 11+ model=self.engine, 12+ messages=[{"role": "user", "content": prompt}], 13+ max_completion_tokens=150, 14 ) 15- raw_text = response.choices[0].text.strip() 16+ raw_text = response.choices[0].message.content.strip() 17 18- answer = response.choices[0].text.strip().lower() 19+ answer = response.choices[0].message.content.strip().lower() 20 21- generated_text = response.choices[0].text.strip() 22+ generated_text = response.choices[0].message.content.strip()
Figure 11: Case study on reproducing a paper, called Self-Instruct: Aligning Language Models with Self-Generated Instructions (Wang et al., 2023) (repository: self-instruct), displaying the files modified during manual debugging.
geval 1# config.yaml 2- name: text-davinci-003 3+ name: gpt-4.1-nano 4 5+ summ_eval_name: data/summeval_1.csv 6+ dialogue_name: data/topical_chat.csv 7+ hallucination_name: data/qags.csv 1# config.py 2- "name": "gpt-4", 3+ "name": "gpt-4o-mini", 1# llm_evaluator.py 2+ from openai import OpenAI 3+ client = OpenAI(api_key=os.environ.get("OPENAI_API_KEY")) 4 5- self.samples: int = int(params.get("samples", 20)) if self.model_name.lower() == "gpt-4" else 1 6+ self.samples: int = int(params.get("samples", 20)) if "gpt-4" in self.model_name.lower() else 1 7 8- if self.model_name.lower() == "gpt-4": 9- response = openai.ChatCompletion.create( 10+ if "gpt-4" in self.model_name.lower(): 11+ response = client.chat.completions.create( 12 13- max_tokens=self.max_tokens, 14+ max_completion_tokens=self.max_tokens, 15 16- text: str = response["choices"][0]["message"]["content"].strip() 17+ text = response.choices[0].message.content.strip() 18 19- choice["message"]["content"].strip() for choice in response.get("choices", []) 20+ choice.message.content.strip() for choice in response.choices 21 22- if score_normalization and self.model_name.lower() == "gpt-4": 23+ if score_normalization and "gpt-4" in self.model_name.lower():
Figure 12: Case study on reproducing a paper, called G-Eval: NLG Evaluation using GPT-4 with Better Human Alignment(Liu et al., 2023) (repository: geval), displaying the files modified during manual debugging.